5 years ago, Adam was half as old as he will be 22 years from now.
How old is Adam now?

Answer by ikleyn(52795)

.

    2*(a-5) = a + 22


    2a - 10 = a + 22

    2a - a = 22 + 10

       a   = 32.


ANSWER.  ADAM is 32 years old now.

Solved.